# An Embedding is Worth a Thousand Noisy Labels

### v1

@@ 4 Experiments and results
Initially, in Section 4.1, we justify the choice of the DINOv2 Large backbone [34]. Next, in Sections 4.2 – 4.5, we evaluate WANN against robust loss functions on noisy real-world, limited, and medical data. In Section 4.6, we assess the robustness of the weighted adaptive neighborhood in the context of heavily imbalanced datasets. We further demonstrate the effectiveness of the proposed Filtered LDA (FLDA) approach in Section 4.7. Finally, in Section 4.8, we qualitatively explore the explainability benefits of our approach on noisy datasets. It is important to note that none of the datasets used in our quantitative experiments were part of DINOv2’s training data, as detailed in Table 15 of [34].
@@ 4.1 Backbone
In order to achieve satisfactory $k$-NN performance, a high-quality feature representation is essential to ensure close proximity of similar objects. To evaluate this, we initially compare the performance of ResNet50111resnet50.tv_in1k and ResNet101222resnet101.tv_in1k [20], both pre-trained on ImageNet-1k. In the realm of Vision Transformers, we explore self-supervised pre-training with MAE [18] (Base333vit_base_patch16.mae and Large444vit_large_patch16.mae), CLIP pre-training [38] (Base555vit_base_patch16_clip_224.openai and Large666vit_large_patch14_clip_336.openai), and lastly DINOv2 pre-training [34] (Base777vit_base_patch14_dinov2.lvd142m and Large888vit_large_patch14_dinov2.lvd142m). All pre-trained models are obtained from HuggingFace (timm). We report WANN’s classification accuracy on CIFAR-10 and CIFAR-100 [25].
Results   Table 1 displays the results, clearly confirming DINOv2 as backbone choice. Results further reveal a notable decline in performance for all backbones as the number of classes increases. This trend is expected due to increasing overlaps of class-specific feature distributions within the embedding space. Additionally, obtaining ground truth labels for a larger set of visually more ambiguous classes poses more challenges associated with label quality. Notably, comparing MAE ViTs with ResNets, both pre-trained on ImageNet, it is possible to observe a substantially lower performance of MAE ViTs. Despite the higher number of parameters, the self-supervised pre-training paradigm appears to be not beneficial in this context. Furthermore, the discrepancy between CLIP and DINOv2 can be attributed to the divergent pre-training procedures: CLIP is trained on pairs of image-text, whereas DINOv2 exclusively relies on image data. Beyond computational efficiency, the DINOv2 framework allows to disregard potential ambiguities introduced by textual descriptions. For this reason, it also required considerably less data, as DINOv2 was trained on $142$ million images while CLIP on $400$ millions of image-text pairs. Thus, in scenarios where text embeddings are not explicitly needed, as in our case, opting for an image-centric training paradigm seems beneficial.
Backbone #Par #Dim C-10 C-100
@@
Table 1: WANN’s classification performance (accuracy) over eight different feature spaces, on CIFAR-10 and CIFAR-100, which contains 10 and 100 classes, respectively. Note that “#Par” denotes the number of parameters and “#Dim” denotes the feature dimension.
@@
Table 2: Accuracy ($\uparrow$) on two clean and noisy datasets [49]. The two best results are highlighted in bold. Note that NR stands for Noise Rate. CIFAR-10N includes three annotations per image, and Noisy split denotes various aggregation strategies. Following the proposed naming convention, Aggr refers to a majority voting, R-$i$ ($i\in\{1,2,3\}$) denotes the $i$-th submitted label for each image, and Worst denotes the selection of only wrong labels, if applicable. Conversely, CIFAR-100N presents just one annotation per image, thereby offering one single noisy split.
4.2 Real-world noisy labels
We compared ANN and WANN with lightweight techniques falling into the category of robust loss functions. As commonly employed for assessing the feature space quality [34], we trained a single linear layer over the image embeddings. The embeddings were normalized using training set statistics, excluding 15% for validation in linear training. The linear probing utilizes the Adam optimizer with a learning rate of $1{\times}10^{-4}$ for $100$ epochs, with early stopping after 15 epochs. These settings were consistently applied to all experiments in this manuscript. Following the publicly available losses and hyperparameters proposed by [54], we compared: Cross Entropy (CE) as a baseline, Focal Loss (FL), Mean Absolute Error (MAE), Generalized Cross Entropy (GCE), Symmetric Cross Entropy (SCE), Active Passive Losses (NCE+MAE, NCE+RCE, NFL+RCE), Asymmetric Losses (NCE+AGCE, NCE+AUL, NCE+AEL), and Active Negative Losses (ANL+CE, ANL+FL). Additionally, we include Early Learning Regularization (ELR), using the suggested hyperparameters. Since the embeddings are pre-generated, no data augmentation was applied. Both ANN and WANN use $(k_{\text{min}},k_{\text{max}})=(11,51)$. We measure the accuracy on two publicly available datasets, namely CIFAR-10N and CIFAR-100N [49], which are “noisy versions” of the established CIFAR datasets. CIFAR-10N contains three human annotations per image, proposing different noisy training splits with varying levels of noise rates (NR). Conversely, CIFAR-100N presents solely one annotation per image, thereby offering one single noisy split. Due to the potential sensitivity of linear training to the choice of the random seed, we show the accuracy as the mean and standard deviation computed over five runs.
Results   Table 2 shows the results of the designed experiment for real-world noisy labels. The introduced weighting scheme in WANN substantially enhances robustness compared to ANN, particularly in high-noise conditions. Additionally, WANN exhibits greater advantages compared to robust loss functions in scenarios with limited noise ratios (${<}20\%$), showing a notable lower performance drop than robust loss functions on CIFAR-10N Aggr and R1, respectively. Active Negative Losses (ANL-CE and ANL-FE) rank among the top robust loss functions, approaching WANN’s performance. Furthermore, while NCE-AGCE exhibits a higher accuracy on CIFAR-100N compared to WANN, it notably presents a higher drop in performance with respect to the clean performance.
@@ 4.2 Real-world noisy labels
Table 3: Accuracy ($\uparrow$) on limited noisy data settings. Both experiments are conducted on CIFAR-10, with $50$ noisy samples per class (left) and $100$ noisy samples per class (right). We highlight the significantly best method (paired $t$-test, $p<0.05$) or the top two methods if their difference is not statistically significant.
4.3 Limited noisy data
The inherent limitation of deep models lies in their reliance on a large amount of clean annotated data. In contrast, $k$-NN models demand less data, exhibiting potential robustness, particularly in the presence of label noise. To demonstrate this, two subsets are randomly subsampled from CIFAR-10, each containing only $50$ and $100$ samples per class. This experiment compares robust loss functions against ANN and WANN. As we do require a validation set for linear probing, we further extract $15\%$ of noisy samples to use as validation set, following previous experiments.
Artificial noise is introduced in CIFAR-10 using traditional methods: symmetric [35], asymmetric [39], and instance-dependent [50] noise. Symmetric noise involves the random switching of one label with any other label in the dataset. Asymmetric noise flips a given label with the adjacent one (i.e., label $i\mapsto i+1\mod C$, where $C$ is the number of classes). Finally, instance-dependent noise takes into account the relationships between the observations within the dataset. Following [58], we report the accuracy on 60% symmetric noise, 30% asymmetric noise, and 40% instance-dependent noise. Due to the stochasticity in subsampling and noise generation, we average the results across five seed runs.
Results   As shown in Table 3, across all experiments (i.e., two sample sizes and three noise injections each), WANN is clearly superior to both ANN and robust loss functions. The advantage of adaptive methods over robust loss functions is particularly evident under heavy noise conditions ($40\%$ inst. and $60\%$ asym.). Although this performance gap narrows slightly as noise rate decreases and data size increases (i.e., under less challenging conditions), the weighting scheme introduced in WANN continues to demonstrate clear benefits over ANN with minimal overhead. Moreover, a paired $t$-test ($p{<}0.05$) conducted across five individual runs confirms that WANN is the significantly best method under heavy noise conditions, underscoring its effectiveness in challenging scenarios. Figure 3 illustrates that under asymmetric noise, incorrect labels may cluster closely in the embedding space. This in fact poses challenges for the traditional $k$-NN but is mitigated by WANN. Notably, with 100 samples per class and $30\%$ asymmetric noise, while NCE-MAE approaches the performance of ANN, it does not surpass WANN. Additionally, while the difference between WANN and NCE-MAE is not statistically significant, WANN outperforms NCE-MAE by $1.37\%$ with an approximately four times lower standard deviation, suggesting potentially greater generalization capabilities.
Figure 3: t-SNE projection of a CIFAR-10 subset and its noisy (20% asymmetric) counterpart.
@@ 4.3 Limited noisy data
4.4 Limited real-world noisy data
We further assess classification performance on five stratified subsets of Animal-10N [41] consisting of $\{100,500,1000,2500,5000\}$ total samples. It is a widely adopted dataset for label noise benchmarks [54] and includes $50{,}000$ samples across $10$ classes, with approximately  ${\approx}\,8\%$ noisy labels. We compare the performance of ANN and WANN against three representative loss functions (cf. Figure 4). For linear probing, $15\%$ of the samples are retained as a validation set. No loss-hyperparameters were tuned for these datasets; instead, we used those found for CIFAR-10. For adaptive methods we adjust $(k_{\text{min}},k_{\text{max}})=(5,11)$ solely on the smallest sample size, as we expect  ${\approx}\,10$ samples per class. Accuracy is averaged across five runs.
Results   As shown in Figure 4, while linear probing approaches $k$-NN methods (i.e., ANN and WANN) with larger subsets, its effectiveness diminishes with limited data. Specifically, when reduced to about $10$ (noisy) samples per class, linear methods substantially underperform against WANN and ANN, exhibiting a much higher variance. This performance gap narrows only when the sample size is increased to about $2{,}500$ samples.
@@ 4.5 Generalization to medical data
To demonstrate the generalization capabilities and effectiveness of WANN, we provide another pool of experiments on two challenging medical datasets, which are semantically far from the pre-training data of the employed foundation model. Specifically, we used two datasets from the MedMNIST+ dataset collection [55], namely BreastMNIST and DermaMNIST, each with resolution $224{\times}224$. BreastMNIST is a binary dataset with only $546$ training samples. We limit our noise injection to symmetric noise at $20\%$, $30\%$, and $40\%$ noise ratios. In fact, on a binary dataset, symmetric and asymmetric noises are equal. DermaMNIST, on the other hand, is a multi-class dataset with $7$ classes and $7{,}007$ training samples. Thus, we apply the same previous noise settings (cf. Table 4). As before, we used the loss-hyperparameters defined for CIFAR-10. We report the average accuracy over five runs.
Results   Table 4 presents the results for the medical datasets. Notably, on BreastMNIST, WANN exhibits the highest accuracy with both clean and noisy training sets, highlighting its effectiveness in scenarios with limited data. Additionally, it significantly outperforms (paired $t$-test, $p<0.05$) ANN and robust loss functions under severe label noise (i.e., $40\%$ symmetric noise). While it might be expected that $k$-NN-based approaches would outperform linear methods on a binary problem with a small dataset, the weighting scheme in WANN demonstrates clear benefits on DermaMNIST as well. In fact, although linear methods outperform ANN, they do not surpass WANN, which consistently achieves a higher accuracy and lower standard deviation, demonstrating greater robustness.
@@
Table 4: Accuracy ($\uparrow$) on BreastMNIST and DermaMNIST. We highlight the significantly best method (paired $t$-test, $p<0.05$) or the top two methods if their difference is not statistically significant.
@@ 4.6 Long-tailed noisy data
While $k$-NN is inherently robust, the selection of an appropriate value of $k$ is critical and not trivial, especially considering the unknown noise rate and potential class imbalances. In this experiment, we assess the performance of ANN and WANN, as compared to two fixed $k$-NN approaches. As discussed in the method section, both adaptive methods used $(k_{\text{min}},k_{\text{max}})=(11,51)$. They are compared with their respective fixed $k$-NN counterparts, namely $11$-NN and $51$-NN. For benchmarking on long-tailed problems, we use CIFAR-10LT and CIFAR-100LT, following prior studies [10, 14] with imbalance ratios of $1\%$ and $10\%$, respectively. It denotes the ratio between the least and most frequent class, with an exponentially decaying imbalance across all other classes. In CIFAR-10LT, the majority class has $5000$ samples, while the minority class consists of $50$ samples. Consequently, in CIFAR-100LT, we have $500$ and $50$ samples, respectively. As in previous experiments, noise is artificially injected. However, here we emphasize the generalizability across different noise patterns and severities. Thus, we inject symmetric noise ranging from $20\%$ to $60\%$, asymmetric noise from $20\%$ to $40\%$, and instance-dependent noise from $20\%$ to $40\%$, reporting the accuracy over five runs.
Results   The results presented in Figure 5 align with our expectations. Under limited noise conditions, a lower value of $k$ ($k=11$) generally achieves higher performance, while a higher value ($k=51$) shows greater robustness against more severe noise levels. In contrast, adaptive methods consistently approach the best performances across various noise rates and patterns. WANN’s weighting scheme consistently exhibits higher gains compared to ANN, especially under heavy asymmetric and instance-dependent noise. Additionally, considering all experiments and respective seed runs (due to the dataset variations from the stochastic noise injection), WANN emerges as the significantly best method, confirmed by the Wilcoxon signed-rank test ($p<0.05$).
@@
Figure 5: Accuracy of fixed and adaptive $k$-NN based approaches on CIFAR-10LT and CIFAR-100LT, with $1\%$ and $10\%$ imbalance ratios. The imbalance ratio denotes the ratio between the least and the most frequent class, with an exponentially decaying imbalance between all other classes. Notably, WANN is close to the best performance across any noise pattern and severity, confirming its robustness and generalizability. Furthermore, WANN is the significantly best method across all experiments and datasets (Wilcoxon signed-rank test, $p<0.05$).

### v3

@@ 4 Experiments and results
Initially, in Section 4.1, we justify the choice of the DINOv2 Large backbone (Oquab et al., 2023). Next, in Sections 4.2 – 4.5, we evaluate WANN against robust loss functions on noisy real-world, limited, and medical data. In Section 4.6, we assess the robustness of the weighted adaptive neighborhood in the context of heavily imbalanced datasets. We further demonstrate the effectiveness of the proposed Filtered LDA (FLDA) approach in Section 4.7. Moreover, in Section 4.8, we qualitatively explore the explainability benefits of our approach on noisy datasets. It is important to note that none of the datasets used in our quantitative experiments were part of DINOv2’s pretraining data, as detailed in Table 15 of (Oquab et al., 2023). Finally, Appendix AB demonstrate the generalizability of WANN utilizing an additional backbone and its robustness with respect to the proposed reliability score, respectively.
@@ 4.1 Backbone
In order to achieve satisfactory $k$-NN performance, a high-quality feature representation is essential to ensure close proximity of similar objects. To evaluate this, we initially compare the performance of ResNet50 and ResNet101 (He et al., 2016), both pre-trained on ImageNet-1k. In the realm of Vision Transformers, we explore self-supervised pre-training with MAE (He et al., 2022) (Base and Large), CLIP pre-training Radford et al. (2021) (Base and Large), and lastly DINOv2 pre-training (Oquab et al., 2023) (Base and Large). All pre-trained models are obtained from HuggingFace (timm). We report WANN’s classification accuracy on CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), utilizing $(k_{\text{min}},k_{\text{max}})=(11,51)$.
Results   Table 1 displays the results, clearly confirming DINOv2 as backbone choice. Results further reveal a notable decline in performance for all backbones as the number of classes increases. This trend is expected due to increasing overlap of class-specific feature distributions within the embedding space. Additionally, obtaining ground truth labels for a larger set of visually more ambiguous classes poses more challenges associated with label quality. Notably, comparing MAE ViTs with ResNets, both pre-trained on ImageNet, it is possible to observe a substantially lower performance of MAE ViTs. Despite the higher number of parameters, the self-supervised pre-training paradigm appears to be not beneficial in this context. Furthermore, the discrepancy between CLIP and DINOv2 can be attributed to the divergent pre-training strategies: CLIP is trained on pairs of image-text, whereas DINOv2 exclusively relies on image data. Beyond computational efficiency, the DINOv2 framework allows to disregard potential ambiguities introduced by textual descriptions. For this reason, it also required considerably less data, as DINOv2 was trained on $142$ millions of images while CLIP on $400$ millions of image-text pairs. Thus, in scenarios where text embeddings are not explicitly needed, as in our case, opting for an image-centric training paradigm is beneficial.
Backbone #Par #Dim CIFAR-10 CIFAR-100
@@
Table 1: WANN’s classification performance (accuracy $\uparrow$) over eight different feature spaces, on CIFAR-10 and CIFAR-100, which contains 10 and 100 classes, respectively. Note that “#Par” denotes the number of parameters and “#Dim” denotes the feature dimension.
4.2 Real-world noisy labels
We compared ANN and WANN with lightweight techniques falling into the category of robust loss functions. As commonly employed for assessing the feature space quality (Oquab et al., 2023), we trained a single linear layer over the image embeddings. The embeddings were normalized using training set statistics, excluding 15% for validation in linear training. The linear probing utilizes the Adam optimizer with a learning rate of $1{\times}10^{-4}$ for $100$ epochs, with early stopping after 5 epochs. These settings were consistently applied to all experiments in this manuscript. Following the publicly available losses and hyperparameters proposed by Ye et al. (2023), we compared: Cross Entropy (CE) as a baseline, Focal Loss (FL), Mean Absolute Error (MAE), Generalized Cross Entropy (GCE), Symmetric Cross Entropy (SCE), Active Passive Losses (NCE-MAE, NCE-RCE, NFL-RCE), Asymmetric Losses (NCE-AGCE, NCE-AUL, NCE-AEL), and Active Negative Losses (ANL-FL,ANL-CE-ER). Additionally, we include Early Learning Regularization (ELR) (Liu et al., 2020), a widely adopted baseline related to label correction. Since the embeddings are pre-generated, no data augmentation was applied. Both ANN and WANN use $(k_{\text{min}},k_{\text{max}})=(11,51)$. We measure the accuracy on two publicly available datasets, namely CIFAR-10N and CIFAR-100N (Wei et al., 2022), which are “noisy versions” of the established CIFAR datasets. CIFAR-10N contains three human annotations per image, proposing different noisy training splits with varying levels of noise rates (NR). Conversely, CIFAR-100N presents solely one annotation per image, thereby offering one single noisy split. Due to the potential sensitivity of linear training to the choice of the random seed, we show the accuracy as the mean and standard deviation computed over five runs.
Results   Table 2 shows the results of the designed experiment for real-world noisy labels. The introduced weighting scheme in WANN substantially enhances robustness compared to ANN, particularly in high-noise conditions. Additionally, WANN exhibits greater advantages compared to robust loss functions in scenarios with limited noise ratios ($<20\%$), showing a notable lower performance drop than robust loss functions on CIFAR-10N Aggr and R1, respectively. Active Negative Losses (ANL-CE-ER and ANL-FE) rank among the top robust loss functions, approaching WANN’s performance. Furthermore, while NCE-AGCE exhibits a higher accuracy on CIFAR-100N compared to WANN, it notably presents a higher performance drop with respect to the clean data ($7.15\%$ versus $6.24\%$), indicating lower robustness.
@@ 4.2 Real-world noisy labels
Table 2: Accuracy ($\uparrow$) on two clean and noisy datasets (Wei et al., 2022). The two best results are highlighted in bold. Note that NR stands for Noise Rate. CIFAR-10N includes three annotations per image, and Noisy split denotes various aggregation strategies. Following the proposed naming convention, Aggr refers to a majority voting, R-$i$ ($i\in\{1,2,3\}$) denotes the $i$-th submitted label for each image, and Worst denotes the selection of only wrong labels, if applicable. Conversely, CIFAR-100N presents just one annotation per image, thereby offering one single noisy split.
4.3 Limited noisy data
The inherent limitation of deep models lies in their reliance on a large amount of clean annotated data. In contrast, $k$-NN models demand less data, exhibiting potential robustness, particularly in the presence of label noise. To demonstrate this, two subsets are randomly sampled from CIFAR-10, each containing only $50$ and $100$ samples per class. This experiment compares robust loss functions against ANN and WANN.
Due to the limited dataset size, we further extract an additional noisy validation set of $250$ samples (i.e., $25$ samples per class). However, WANN does not require this additional data, which is often difficult to obtain in limited-data settings.
Artificial noise is introduced in CIFAR-10 using traditional methods: symmetric (Patrini et al., 2017), asymmetric (Scott et al., 2013), and instance-dependent (Xia et al., 2020) noise. Symmetric noise involves the random switching of one label with any other label in the dataset. Asymmetric noise flips a specific label to another fixed, incorrect label (e.g., “deer” mislabeled as “horse”), simulating systematic class confusion. Finally, instance-dependent noise takes into account the relationships between the features within the dataset. Following Zhu et al. (2022), we report the accuracy on 60% symmetric noise, 30% asymmetric noise, and 40% instance-dependent noise. Following Ye et al. (2023), our asymmetric flips are defined as: TRUCK $\rightarrow$ AUTOMOBILE, BIRD $\rightarrow$ AIRPLANE, DEER $\rightarrow$ HORSE, CAT $\leftrightarrow$ DOG. Due to the stochasticity in subsampling and noise generation, we average the results across five seed runs.
Results   As shown in Table 3, across all experiments (i.e., two sample sizes and three noise injections each), WANN is clearly superior to both ANN and robust loss functions. The advantage of adaptive methods over robust loss functions is particularly evident under heavy noise conditions (e.g., $60\%$ sym.). Although this performance gap narrows slightly as data size increases (i.e., under less challenging conditions), the weighting scheme introduced in WANN continues to demonstrate clear benefits over ANN with minimal overhead. Moreover, a paired $t$-test ($p<0.05$) conducted across five seed runs confirms that WANN is the significantly best method in all but two cases (cf. Table 3). Furthermore, with $100$ samples per class, WANN consistently exhibits lower standard deviation across seed runs, thereby confirming its higher robustness against competing methods. Figure 3 illustrates that under asymmetric noise, incorrect labels may cluster closely in the embedding space. This in fact poses challenges for the traditional $k$-NN but is mitigated by WANN. Notably, while linear methods approach and surpass the performance of ANN with $100$ samples per class, they do not outperform WANN.
@@ 4.3 Limited noisy data
Table 3: Accuracy ($\uparrow$) on limited noisy data settings. Both experiments are conducted on CIFAR-10, with $50$ noisy samples per class (left) and $100$ noisy samples per class (right). We bold the top two methods and underline the significantly best one if their difference is statistically significant (paired $t$-test, $p<0.05$).
4.4 Limited real-world noisy data
We further assess classification performance on four stratified subsets of Animal-10N (Song et al., 2019) consisting of $\{500,1000,2500,5000\}$ total samples. It is a widely adopted dataset for label noise benchmarks (Ye et al., 2023) and includes $50{,}000$ samples across $10$ classes, with approximately  ${\approx}\,8\%$ noisy labels. We compare the performance of ANN and WANN against three representative loss functions (cf. Figure 4). As in Section 4.3, solely for linear probing methods, we extract an additional noisy validation set of 250 samples. Furthermore, we do not tune loss-specific hyperparameters for this dataset; instead, we used those found for CIFAR-10. We report the results obtained across five seed runs.
Results   As shown in Figure 4, while linear probing approaches $k$-NN methods (i.e., ANN and WANN) with larger subsets, its effectiveness diminishes with limited data. Specifically, when reduced to about $50$ (noisy) samples per class, linear methods substantially underperform against ANN and WANN, exhibiting a larger $95\%$ confidence interval. This performance gap narrows only when the sample size is increased to about $2{,}500$ and $5{,}000$ samples.
Figure 3: t-SNE projection of a CIFAR-10 subset and its noisy (30% asymmetric) counterpart.
@@ 4.5 Generalization to medical data
To demonstrate the generalization capabilities and effectiveness of WANN, we provide another pool of experiments on two challenging medical datasets, which are semantically far from the pre-training data of the employed foundation model. Specifically, we used two datasets from the MedMNIST+ dataset collection (Yang et al., 2023), namely BreastMNIST and DermaMNIST, each with resolution $224{\times}224$. BreastMNIST is a binary dataset with only $546$ training samples. Given its binary nature, we limit our noise injection to symmetric noise at $20\%$, $30\%$, and $40\%$ noise ratios. DermaMNIST, on the other hand, is a multi-class dataset with $7$ classes and $7{,}007$ training samples. Thus, we apply the same noise settings as previously (cf. Table 4). For asymmetric noise, we shift each class by one in a circular manner. We use the loss-hyperparameters defined for CIFAR-10 and report average accuracy over five runs.
Results   Table 4 presents the results for the medical datasets. Notably, on BreastMNIST, WANN and ANN exhibit the highest accuracy with both clean and noisy training sets. Furthermore, WANN outperforms ANN with increased symmetric noise. While it might be expected that $k$-NN-based approaches would outperform linear methods on a binary problem with a small dataset, the weighting scheme in WANN demonstrates clear benefits on DermaMNIST as well, with a significant improvement on instance dependent label noise. In fact, although GCE outperforms ANN with symmetric noise, it does not surpass WANN, which consistently achieves higher accuracy and lower standard deviation, demonstrating greater robustness.
@@
Table 4: Accuracy ($\uparrow$) on BreastMNIST and DermaMNIST. We bold the top two methods and underline the significantly best one if their difference is statistically significant (paired $t$-test, $p<0.05$).
@@ 4.6 Long-tailed noisy data
While $k$-NN is inherently robust, the selection of an appropriate value of $k$ is critical and not trivial, especially considering the unknown noise rate and potential class imbalances. In this experiment, we assess the performance of ANN and WANN, as compared to two fixed $k$-NN approaches. As discussed in the method section, both adaptive methods used $(k_{\text{min}},k_{\text{max}})=(11,51)$. They are compared with their respective fixed $k$-NN counterparts, namely $11$-NN and $51$-NN. For benchmarking on long-tailed problems, we use CIFAR-10LT and CIFAR-100LT, following prior studies (Cao et al., 2019; Du et al., 2023), with imbalance ratios of $1\%$ and $10\%$, respectively. It denotes the ratio between the least and most frequent class, with an exponentially decaying imbalance across all other classes. In CIFAR-10LT, the majority class has $5000$ samples, while the minority class consists of $50$ samples. Consequently, in CIFAR-100LT, we have $500$ and $50$ samples, respectively.
As in previous experiments, noise is artificially injected. To emphasize the generalizability across different noise patterns and severities, we inject symmetric noise ranging from $20\%$ to $60\%$, asymmetric noise from $20\%$ to $40\%$, and instance-dependent noise from $20\%$ to $40\%$, reporting the accuracy over five runs.
Results   The results presented in Figure 5 align with our expectations. Under limited noise conditions, a lower value of $k$ ($k=11$) generally achieves higher performance, while a higher value ($k=51$) shows greater robustness against more severe noise levels. In contrast, adaptive methods consistently approach the best performances across various noise rates and patterns. WANN’s weighting scheme exhibits higher gains compared to ANN, especially under heavy asymmetric and instance-dependent noise. Additionally, considering all experiments and respective seed runs (due to the dataset variations from the stochastic noise injection), WANN emerges as the significantly best method (Wilcoxon signed-rank test, $p<0.05$).
@@
Figure 5: Accuracy of fixed and adaptive $k$-NN based approaches on CIFAR-10LT and CIFAR-100LT, with $1\%$ and $10\%$ imbalance ratios. The imbalance ratio denotes the ratio between the least and the most frequent class, with an exponentially decaying imbalance between all other classes. Notably, WANN is close to the best performance across any noise pattern and severity. Furthermore, WANN is the significantly best method across all experiments and datasets (Wilcoxon signed-rank test, $p<0.05$).